Shift report 20030517 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat May 17 2003 - 16:05:25 EDT

  • Next message: Apache: "Shift report 20030518 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat May 17 16:05:24 2003, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#14 pages 112 - 112
    
    Shift Summary:
    ------------------------------------------------------------
    Beam:
    Injecting at ~ 9:21 took about 1 hour to steer (first time since
    star spin rotators were commissioned.
    Rate rather low ~ 11 K at start;
    Beam effectively stop being useful at 13:45 == 3:45 hours.
    
    Data runs:
    FS 4 deg 1/8A 8978,8979  
                    1/1A 8980 
    FS 3 deg  1/1A 8981,8982
    
    MRS 90deg 8979,8980 26K
    MRS 90 deg 350A: 
    
    Magnet PS
    D1 is still going hot and smelly.
    -- record PS temp (go to PS room)
    -- turn down to 0 while not in stored, and cogged mode.
    
    D2 dropped to fault once (not while running)
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat May 17 2003 - 16:10:04 EDT